Meyer & Co. expects its EBIT to be $115,000 every year forever. The firm can borrow at 9 percent. The company currently has no debt, and its cost of equity is 13 percent and the tax rate is 23 percent. The company borrows $168,000 and uses the proceeds to repurchase shares.

  

a.

What is the cost of equity after recapitalization? (Do not round intermediate calculations and enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.)

b. What is the WACC? (Do not round intermediate calculations and enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.)
Answer #1
Solution:
a. Cost of Equity 13.94 %
Working Notes:
Cost of Equity after recapitalization means debt is introduced and the firm become Levered firm and the cost of equity (levered) is asked.
Unlevered value of the firm Vu
Vu = EBIT x (1- tax rate )/ Cost of Equity unlevered
Vu = EBIT x (1- tax rate )/R0
Vu = 115000 x (1- 0.23 )/13%
Vu = 681153.846154
Levered value of the firm VL
VL = Vu + Borrowing x tax rate
VL = 681153.846154 + 168000 x 23%
VL =719793.846154
Now using M&M Proposition II with taxes
value of the levered firm VL = Borrowing + Value of Equity
VL = B + S
719793.846154 = 168000 + S
Market value of Equity S = 719793.846154 - 168000
Market value of Equity S =551793.846154
Using M&M Proposition II with taxes
RS = R0+ (R0-RB)(B/S)(1- Tc)
Where
B/s = debt equity ratio =168000/551793.846154
RS is cost of equity capital levered = ??
R0 is cost of equity capital of unlevered firm = 13%
RB is Cost of Borrowing = 9%
Tc = tax rate = 23%
RS= R0+ (R0-RB)(B/S)(1-Tc)
Rs =13% + (13% - 9%) x (168000/551793.846154) x ( 1- 0.23)
Rs =13% + 0.93774152%
Rs =13.93774152%
Rs =13.94%
b. WACC 12.30 %
Working Notes:
After tax cost of debt (Kd) = Cost of debt = 9% x (1-tax rate)
Kd= =9% x (1-0.23) = 6.93%
Cost of common equity (Ke)=13.93774152%
VL = B + S
Market value of Equity S =551793.846154
VL =719793.846154
B = 168,000
WACC= Ke x S/V + Kd   x B/V
WACC = 13.93774152% x (551793.846154/719793.846154) + 6.93% x (168000/719793.846154)
WACC = 0.123021335
WACC = 12.30%
Please feel free to ask if anything about above solution in comment section of the question.